OpenStudy (anonymous):

3radical2 + 5radical8= ?

OpenStudy (anonymous):

\[3\sqrt{2}+5\sqrt{8}\] One way I can see this is that, convert radical 2 and 8 into decimal, then multiply: 4.24+14.14 18.38

OpenStudy (anonymous):

Can you simplify 5radical8 to something with some number * radical 2?

OpenStudy (anonymous):

does it help to know that \[\sqrt{8}=\sqrt{4\times 2}=\sqrt{4}\times \sqrt{2}=2\sqrt{2}\]?

OpenStudy (anonymous):

so you have \[3\sqrt{2}+10\sqrt{2}\] now combine like terms

OpenStudy (anonymous):

yes it does thanks